# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v2

@@ Abstract
Reconstruction attacks and defenses are essential in understanding the data leakage problem in machine learning. However, prior work has centered around empirical observations of gradient inversion attacks, lacks theoretical justifications, and cannot disentangle the usefulness of defending methods from the computational limitation of attacking methods. In this work, we propose to view the problem as an inverse problem, enabling us to theoretically, quantitatively, and systematically evaluate the data reconstruction problem. On various defense methods, we derived the algorithmic upper bound and the matching (in feature dimension and model width) information-theoretical lower bound on the reconstruction error for two-layer neural networks. To complement the theoretical results and investigate the utility-privacy trade-off, we defined a natural evaluation metric of the defense methods with similar utility loss among the strongest attacks. We further propose a strong reconstruction attack that helps update some previous understanding of the strength of defense methods under our proposed evaluation metric.
@@ 1 Introduction
Machine learning research has transformed the technical landscape across various domains but also raises privacy concerns potentially (Li et al., 2023; Papernot et al., 2016). Federated learning (Konečný et al., 2016; McMahan et al., 2017), a collaborative multi-site training framework, aims to uphold user privacy by keeping user data local while only exchanging model parameters and updates between a central server and edge users. However, recent studies on reconstruction attacks (Yin et al., 2021; Huang et al., 2021) highlight privacy vulnerabilities even within federated learning. Attackers can eavesdrop on shared trained models and gradient information and reconstruct training data using them. Even worse, honest but curious servers can inadvertently expose training data by querying designed model parameters. Some defending methods are also proposed and analyzed (Zhu et al., 2019; Wei et al., 2020b; Zhang et al., 2018).
However, previous research on reconstruction attacks and defenses centered around empirical studies and lacked theoretical guarantees. Moreover, those experiments over defenses usually focus only on one specific attack, making the evaluation of defenses’ strength untrustworthy. Empirical understanding of attacking and defending methods is affected by factors like optimization challenges of non-convex objectives. Under the heuristic attacking methods from prior work, it is hard to disentangle the following two possibilities: 1) the algorithmic or computational barrier of the specific attack and 2) the defense’s success. Thus, a systematical evaluation of defenses involving information theoretical analysis and eliminating the attack algorithm’s computational barrier is needed.
A line of theoretical work analyzing privacy attacks is differential privacy (DP) (Dwork, 2006), which guarantees privacy in data re-identification or membership inference. However, there are scenarios when data identity information is not sensitive, but data itself is, and DP is unsuitable for interpreting data reconstruction attacks (Guo et al., 2022; Hayes et al., 2024). Although some previous works discussed the reconstruction attack with the more tailored Renyi-DP (Guo et al., 2022; Stock et al., 2022), their setting is when all samples but one are known, and the attacker’s goal is to reconstruct the last data point, which doesn’t apply to the federated learning setting. We will discuss how (Renyi-)DP is too stringent for the data leakage problem under federated learning in Section 2. Some underlying structures of the observation can yield no (Renyi-)DP guarantee but prevent the attacker from reconstructing the data.
@@
In our work, we instead propose a theoretical framework that views the data reconstruction as an inverse problem (Fig. 1): from the gradient $G$ that is generated from the unknown data $S$ and a known function (training loss’s gradient), the adversary intends to reconstruct the input data, and a defender $D$ is to prevent this from perturbing the observations. Our theoretical analysis focuses on 2-layer fully connected networks and gives both upper and lower bounds on the error of this inverse problem under various defenses. The algorithmic upper bounds are based on a specific attack algorithm proposed by (Wang et al., 2023), and the lower bounds are derived from a statistical point of view. To generalize the analysis to real-world models and study the utility-privacy trade-offs, we empirically conduct a comprehensive evaluation of different defenses’ performances. Our comparisons are based on the strongest attack corresponding to each defense, so the result is general instead of especially effective for a specific attack.
@@ 2 Preliminary
In reconstruction attacks, we observe the model iterations in training and recover data from the gradient $\nabla L(\Theta;S)$, where $L$ is the loss function, $\Theta$ is the model and $S$ is an unknown dataset. Then, it can be regarded as an inverse problem to identify unknown signal $S$ from observation $G=\nabla L(\Theta;S)$, where $\nabla L$ is the forward function. We denote the data space by $\mathcal{X}$ and the batch size of data by $B$. Then $S\subset\mathcal{X}^{B}$. Under the inverse problem framework, the effectiveness of attacks and defenses becomes the feasibility and hardness of recovering $S$. Formally, we define a minimax risk of reconstruction error:
@@
Our studies are meant to complement the esteemed DP (Dwork, 2006) to analyze the reconstruction as DP is too strong and sometimes unnecessary for our studied settings. On one hand, scenarios exist when reconstruction is impossible, even when DP does not holdFor instance, one can easily verify that with linear net and $l_{2}$ loss, the observation becomes $\sum_{i=1}^{B}x_{i}$, where no DP is satisfied but individual recovery of $x_{i}$ is impossible for $B>1$ unless specific prior knowledge is assumed.. In addition, the price of achieving DP in the data reconstruction setting is high. For DP-SGD (Abadi et al., 2016), the well-established Gaussian mechanism requires a large variance that will destroy the utility of gradient information:
@@ 3 Theoretical Analysis
In theoretical analysis, we mainly focus on two-layer neural networks $f({\bm{x}};\Theta)=\sum_{j=1}^{m}a_{j}\sigma({\bm{w}}_{j}^{\top}{\bm{x}})$ with $\ell$ be the square loss. Here $a_{j}\sim\mathcal{N}(0,1/m^{2})$ and ${\bm{w}}_{j}\sim\mathcal{N}(0,I)$. For the upper bound, we examine a specific attack method proposed by (Wang et al., 2023). The attack method involves the computation of gradient-related tensor $\sum_{j=1}^{m}g({\bm{w}}_{j})H_{p}({\bm{w}}_{j})$, where $g({\bm{w}}_{j}):=\frac{\partial\ell}{\partial a_{j}}$ and $H_{p}$ is the $p$-th Hermite function. By Stein’s lemma (Stein, 1981; Mamis, 2022), it is approximate to a tensor product $\sum_{i=1}^{B}c_{i}{\bm{x}}_{i}^{\otimes p}$ when ${\bm{w}}_{j}$’s are random Gaussian and we can conduct tensor decomposition to recover data.
@@
For a 2-layer network with noisy gradient, the statistical minimax risk has an upper bound $R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{d}{m}})$ with high probability.
In the following, we analyze the upper and lower bound for the reconstruction under different defenses. The bound is shown in Table 1. Our analysis provides a tight and quantitative analysis of how each defense’s defending strength (such as pruning ratio or noise variance) affects the data reconstruction error (which is inversely proportional to the degree of privacy). In the remaining part of this section, we individually investigate each defense method.
Table 1: The algorithmic upper bound and information-theoretic lower bound of the reconstruction error against different defenses. Some parameters are defined later in the subsection of each defense.
@@ 3.1 Local aggregation
In a realistic setting of federated learning, the observed local update may not be a single gradient. Instead, it can be the result of multiple steps of gradient descent (McMahan et al., 2017). In our analysis, we mainly consider the most simple cases, where local devices train two steps and the observation is $\Theta^{(2)}-\Theta^{(0)}$. Since one step of training does not change the parameters too much, we can approximately assume that the two steps share a same set of parameters. Then the analysis is similar to that without defense.
@@ 3.2 Differential privacy stochastic gradient descent
Differential privacy (Dwork, 2006) algorithms ensure the privacy of data by adding noise to gradient descent steps (Shokri and Shmatikov, 2015; Abadi et al., 2016; Song et al., 2013). In this way, it hides the exact local gradient from the global model. In differential private federated learning (Geyer et al., 2017; Wei et al., 2020a), each client clips the gradient and introduces a random Gaussian noise before updates to the cluster. In the setting of differential privacy stochastic gradient descent (DP-SGD), the update of $a_{j}$ is
@@
Note that under this setting, the information contained in the gradient is disrupted by gradient clipping and random noise. For the upper bound, we first analyze the effect of the two strategies separately and then combine them to bound the error of DP-SGD. If we conduct the tensor-based attack on clipped gradient $\tilde{G}$, there is no defensive effect no matter the clipping threshold and the reconstruction error bound $\tilde{O}(B\sqrt{\frac{d}{m}})$ is the same as that with no defense (even for the hidden terms). For noisy gradient $G+\epsilon_{0}$, we have the error bound $\tilde{O}((B+\sigma_{0})\sqrt{\frac{d}{m}})$, which is the same order to original case if $\sigma_{0}\leq O(\frac{B}{m})$.
@@ 3.3 Secure aggregation
In federated learning, local devices update gradients to the server individually. Secure aggregation (Bonawitz et al., 2016, 2017) is a method that clients $1,\cdots l$ can aggregate their gradients before the cluster, and the cluster can only access to aggregated gradient: $\tilde{G}=\frac{1}{B}\sum_{l=1}^{L}G_{l}B_{l},$ where $G_{l}$ and $B_{l}$ is the gradient update and batch size for $l$-th user respectively and $B=\sum_{l=1}^{L}B_{l}$. In this way, the global server is blocked from knowing each gradient, which prevents privacy leakage. Even though the server can reconstruct the training data, it cannot identify which user has the data. If we only reconstruct data without identifying its source, the aggregated gradient is equivalent to a gradient step with a large batch size $B$, and both bounds remain unchanged.
@@ 3.4 Dropout
Another method to improve the defending effect is dropout (Hinton et al., 2012; Srivastava et al., 2014), a mechanism designed to prevent overfitting. It has been empirically discovered in (Zhu et al., 2019; Huang et al., 2021) that it can defend against reconstruction attack. It randomly drops nodes in a network with a certain probability of $1-p$. In this way, some of the entries in the gradient turn to zero, introducing randomness to the gradient. Thus, it is natural to prevent data leakage with dropout. A two-layer fully connected network with a dropout layer can be seen as a model with an effective width equal to the number of nodes $m^{\prime}\approx(1-p)m$ that has not been dropped. Then $m$ changes into $(1-p)m$ for both upper and lower bounds.
@@ 4.1 Attacks for Defense Methods Evaluation
We apply various defense techniques to the above attack algorithms, and for each defense, we consider the best attack performance across different attack algorithms following Eq. 3.
4.2 Description of the proposed attack algorithm
Regularization on feature matching.
Other regularizations.
Following  (Geiping et al., 2020), we adopt total variation and regularization on batch norm to exploit prior knowledge of natural images and training data. Moreover, inspired by deep image prior (Ulyanov et al., 2018), we employ an untrained Convolutional Neural Network (CNN) to generate an image. This idea is rather similar to (Jeon et al., 2021) where a network is pre-trained for image generation. The final objective of the introduced attack method is
$\arg\min_{x}$ $\mathcal{L}_{\rm grad}(x;\theta,\nabla_{\theta}\mathcal{L}_{\theta}(x^{*},y^{*}))+\alpha_{TV}\mathcal{R}_{\rm TV}+\alpha_{BN}\mathcal{R}_{\rm BN}+\alpha_{f}\mathcal{R}_{\rm feature}(x).$ (4)
5 Experiments
5.1 Experimental setup
Key parameters of defenses.
Key experimental setup for the attack.
We use a subset of 50 CIFAR-10 images to evaluate the attack performance. For all attack algorithms, we assume the batch norm statistics are available following (Geiping et al., 2020). We observe the difficulty of optimization when using cosine-similarity as the gradient inversion loss for the proposed attack algorithm and therefore address this issue by reweighting the gradients by their $\ell_{0}$ norm. See details in the Appendix FGeiping et al., 2020; Zhu et al., 2019 have shown that small batch size is vital for the success of the attack. We intentionally evaluate the attack with two small batch sizes $4$ and $8$ and two small but realistic batch sizes, 16 and 32.
Metrics for reconstruction quality.
5.2 Experiment results
@@
In Section 3, we analyze the performance of different defenses on 2-layer networks. In the experiments, our attack method involves intermediate feature matching, which requires reconstructing feature from a 2-layer network. This setting aligns with our theoretical analysis. Therefore, we can verify the theory by checking the feature reconstruction error in our attack. The results is shown in Table 2, where the reconstruction error increases when the batch size increases for all setting. This matches the theoretical bounds. Moreover, gradient pruning and gradient noise with large scaling has significantly large reconstruction error, which also aligns with our theoretical evaluation.
@@
According to the framework of evaluation, we need to consider strong attacks against various defenses. A key motivation of our proposed attack is its good performance among different defense, which we will verify in the following results. Our method is based on gradient inversion and can be easily added to previous methods (Geiping et al., 2020; Zhu et al., 2019). In Table  3, we compare the state-of-the-art gradient inversion methods with the proposed attacking method, our method outperforms previous methods. We visualize reconstructed images in Figure 2. Without any defense, both (Geiping et al., 2020) and our attack method can recover images well from the gradient, and our method produces higher-quality images. Increasing the batch size makes all methods more difficult to recover well. With defenses, our method shows better robustness – the recovered images are visually more similar to the original image. See Table 7 and Table 8 in Appendix I for summary of reconstruction results.
@@
In our systematic evaluation of different defenses, we measure their strength with the strongest attack: $\mathcal{S}_{D}=\max_{A\in\mathcal{A}}d(S,A(D(G)))$. In order to estimate the maximum reconstruction error, we selected four different attacks to conduct the experiment. The reconstruction RMSE of the four attacks against defenses with various parameters and batch sizes are shown in Table 4. For each setting, we select the attack with the smallest RMSE as the strongest attack. Therefore, among all defenses, gradient pruning with large $p$ and gradient noise with large variance $\sigma_{0}^{2}$ has the best defensive effect.
@@
In order to evaluate the defenses, another key criteria is the utility loss of the defense methods. Most of the defenses prevent models from data leakage by perturbing the gradients in training, which may hurt the training task itself. In our experiment, we measure the utility by the final loss of the training task. We show the final training loss for the defenses in Table 5. Gradient pruning, as one of the defenses having strongest effect towards attacks, causes smaller utility loss than gradient noise. This result indicates that gradient pruning is a better defense than gradient noise. In Figure 3, we plot the relation between the reconstruction error and the utility loss, where each dot represents a method under the setting of a specific batch size and parameter of the defense. With the same level of utility loss, gradient pruning has the largest RMSE, indicating it is the best defense to have $\min_{D\in\mathcal{D}_{\mathcal{U}}}\mathcal{S}_{D}$.

### v3

@@ Abstract
Reconstruction attacks and defenses are essential in understanding the data leakage problem in machine learning. However, prior work has centered around empirical observations of gradient inversion attacks, lacks theoretical grounding, and cannot disentangle the usefulness of defending methods from the computational limitation of attacking methods. In this work, we propose to view the problem as an inverse problem, enabling us to theoretically and systematically evaluate the data reconstruction attack. On various defense methods, we derived the algorithmic upper bound and the matching (in feature dimension and architecture dimension) information-theoretical lower bound on the reconstruction error for two-layer neural networks. To complement the theoretical results and investigate the utility-privacy trade-off, we defined a natural evaluation metric of the defense methods with similar utility loss among the strongest attacks. We further propose a strong reconstruction attack that helps update some previous understanding of the strength of defense methods under our proposed evaluation metric.
@@ 1 Introduction
Machine learning research has transformed the technical landscape across various domains but also raises privacy concerns potentially (Li et al., 2023; Papernot et al., 2016). Federated learning (Konečný et al., 2016; McMahan et al., 2017), a collaborative multi-site training framework, aims to uphold user privacy by keeping user data local while only exchanging model parameters and updates between a central server and edge users. However, recent studies on reconstruction attacks (Yin et al., 2021; Huang et al., 2021) highlight privacy vulnerabilities even within federated learning. Attackers can eavesdrop on shared trained models and gradient information and reconstruct training data using them. Even worse, honest but curious servers can inadvertently expose training data by querying designed model parameters. Some defending methods are also proposed and analyzed.
However, previous research on reconstruction attacks and defenses centered around empirical studies and lacked theoretical guarantees. Moreover, those experiments over defenses usually focus only on one specific attack, making the evaluation of defenses’ strength untrustworthy. Empirical understanding of attacking and defending methods is affected by factors like optimization challenges of non-convex objectives. Under the heuristic attacking methods from prior work, it is hard to disentangle the following two possibilities: 1) the algorithmic or computational barrier of the specific attack and 2) the defense’s success. Thus, a systematical and information-theoretical evaluation of defenses that is independent to the attack method is needed.
A line of theoretical work analyzing privacy attacks is differential privacy (DP) (Dwork, 2006), which guarantees privacy in data re-identification or membership inference. However, there are scenarios when data identity information is not sensitive, but data itself is, and DP is unsuitable for interpreting data reconstruction attacks (Guo et al., 2022; Hayes et al., 2024). An algorithm can yield no DP guarantee but prevent the threat model from data reconstruction (Details deferred to Section 2). Although some previous works discussed the reconstruction attack with more tailored Renyi-DP (Guo et al., 2022; Stock et al., 2022), their setting is when all samples are known, and the attacker’s goal is to reconstruct the last data point, which doesn’t apply to the federated learning setting. We will discuss how (Renyi-)DP is too stringent for the data leakage problem under federated learning in Section 2. Some underlying structures of the observation can yield no (Renyi-)DP guarantee but prevent the attacker from reconstructing the data.
@@
In our work, we propose a framework for evaluating defenses that view the reconstruction attack as an inverse problem (Fig. 1): from the gradient $G$ that is generated from the unknown data $S$ and a known function (objective’s gradient), the adversary intends to reconstruct the input data, and a defender $D$ is to prevent this from perturbing the observations. Our theoretical analysis focuses on 2-layer fully connected networks and gives both upper and lower bounds on the error of this inverse problem under various defenses. We are especially interested in obtaining a quantitative relationship between reconstruction error and the key factors of the learning pipeline, such as data dimension, model architecture (width), and defense strength, which are derived in the form of both algorithmic upper bounds and (matching) information-theoretic lower bound (that is independent of the attack algorithm). To generalize the analysis to real-world models and study the utility-privacy trade-offs, we empirically conduct a comprehensive evaluation of different defenses’ performances. Our comparisons are based on the strongest attack for each defense to exclude the case that some defenses are only effective for a specific attack.
@@ 2 Preliminary
In reconstruction attacks, we observe the model iterations in training and recover data from the gradient $\nabla L(\Theta;S)$, where $L$ is the loss function, $\Theta$ is the model and $S$ is an unknown dataset. Then, it can be regarded as an inverse problem to identify unknown signal $S$ from observation $G=\nabla L(\Theta;S)$, where $\nabla L$ is the forward function. We denote the data space by $\mathcal{X}$ and the batch size of data by $B$. Then $S\in\mathcal{X}^{B}$. Under the inverse problem framework, the effectiveness of attacks and defenses becomes the feasibility and hardness of recovering $S$. Formally, we define a minimax risk of reconstruction error:
@@
Our studies are meant to complement the esteemed DP (Dwork, 2006) to analyze the reconstruction as DP is too strong and sometimes unnecessary for our studied settings. On one hand, scenarios exist when reconstruction is impossible, even when DP does not hold For instance, one can easily verify that with linear net and $l_{2}$ loss, the observation becomes $\sum_{i=1}^{B}x_{i}$, where no DP is satisfied but individual recovery of $x_{i}$ is impossible for $B>1$ unless specific prior knowledge is assumed, for example natural images.. In addition, the price of achieving DP in the data reconstruction setting is high. For DP-SGD (Abadi et al., 2016), the well-established Gaussian mechanism requires a large variance that will destroy the utility of gradient information:
@@
Table 1: The algorithmic upper bound and information-theoretic lower bound of the reconstruction error against different defenses. The bound here is the order with respect to different factors. Some parameters are defined later in the subsection of each defense.
@@ 3 Theoretical Analysis
In theoretical analysis, we mainly focus on two-layer neural networks $f({\bm{x}};\Theta)=\sum_{j=1}^{m}a_{j}\sigma({\bm{w}}_{j}^{\top}{\bm{x}})$ with $\ell$ be the square loss. Here $a_{j}\sim\mathcal{N}(0,1/m^{2})$ and ${\bm{w}}_{j}\sim\mathcal{N}(0,I)$. For the upper bound, we examine a specific attack method proposed by (Wang et al., 2023). The attack method involves the computation of gradient-related tensor $\sum_{j=1}^{m}g({\bm{w}}_{j})H_{p}({\bm{w}}_{j})$, where $g({\bm{w}}_{j}):=\frac{\partial\ell}{\partial a_{j}}$ and $H_{p}$ is the $p$-th Hermite function. By Stein’s lemma (Stein, 1981; Mamis, 2022), it is approximate to a tensor product $\sum_{i=1}^{B}c_{i}{\bm{x}}_{i}^{\otimes p}$ when ${\bm{w}}_{j}$’s are random Gaussian and we can conduct tensor decomposition to recover data, where $B$ is the batch size.
@@
For a 2-layer network with noisy gradient, the statistical minimax risk has a lower bound $R_{L}\geq\tilde{\Omega}(\sigma\sqrt{\frac{d}{m}})$ with high probability.
In the following, we analyze the upper and lower bound for the reconstruction under different defenses. The bound is shown in Table 1. Our analysis provides a tight and quantitative analysis of how each defense’s defending strength (such as pruning ratio or noise variance) affects the data reconstruction error (which is inversely proportional to the degree of privacy).
@@ 3.1 Local aggregation
In a realistic setting of federated learning, the observed local update can be multiple steps of gradient descent (McMahan et al., 2017). In our analysis, we mainly consider the most simple cases, where local devices train two steps and the observation is $\Theta^{(2)}-\Theta^{(0)}$. Since one step of training does not change the parameters too much, we can approximately assume that the two steps share a same set of parameters. Then the analysis is similar to that without defense.
@@ 3.2 Differential privacy stochastic gradient descent
Differential privacy (Dwork, 2006) algorithms hides the exact local gradients by adding noise to gradient descent steps (Shokri and Shmatikov, 2015; Abadi et al., 2016; Song et al., 2013). In differential private federated learning (Geyer et al., 2017; Wei et al., 2020a), each client clips the gradient and introduces a random Gaussian noise before updates to the cluster. In the setting of differential privacy stochastic gradient descent (DP-SGD), the update of $a_{j}$ is
@@
Here $C$ is a constant threshold for gradient clipping. Note that under this setting, the information contained in the gradient is disrupted by gradient clipping and random noise. For the upper bound, the noisy gradient $G+\epsilon_{0}$ will lead to a $\tilde{O}((B+\sigma_{0})\sqrt{\frac{d}{m}})$ error bound, which is the same order to original case if $\sigma_{0}\leq O(\frac{B}{m})$. Gradient clipping will enhance the strength noise though itself has no defensive effect.
@@ 3.3 Secure aggregation
In federated learning, local devices update gradients to the server individually. Secure aggregation (Bonawitz et al., 2016, 2017) is a method that clients can aggregate their gradients before the cluster, and the cluster can only access to aggregated gradient: $\tilde{G}=\frac{1}{B}\sum_{l=1}^{L}G_{l}B_{l},$ where $G_{l}$ and $B_{l}$ is the gradient update and batch size for $l$-th user respectively and $B=\sum_{l=1}^{L}B_{l}$. In this way, the global server is blocked from knowing each gradient, which prevents privacy leakage. Though there is no extra defensive strength compared with directly reconstructing a large batch with size $B$, it cannot identify which user has the data.
@@ 3.4 Dropout
Another method to improve the defending effect is dropout (Hinton et al., 2012; Srivastava et al., 2014), a mechanism designed to prevent overfitting. It has been empirically discovered in (Zhu et al., 2019; Huang et al., 2021) that it can defend against reconstruction attack. It randomly drops nodes in a network with a certain probability of $1-p$. In this way, some of the entries in the gradient turn to zero, introducing randomness to the gradient and prevent data leakage to some degree. A two-layer fully connected network with a dropout layer can be seen as a model with an effective width equal to the number of nodes $m^{\prime}\approx(1-p)m$ that has not been dropped. Then $m$ changes into $(1-p)m$ for both upper and lower bounds.
@@ 4.1 Attacks for Defense Methods Evaluation
We apply various defense techniques to the above attack algorithms, and for each defense, we consider the best attack performance across different attack algorithms following Eq. (3).
4.2 Description of the proposed attack algorithm
Regularization on feature matching.
Other regularizations.
Following  (Geiping et al., 2020), we adopt total variation and regularization on batch norm to exploit prior knowledge of natural images and training data. Moreover, inspired by deep image prior (Ulyanov et al., 2018), we employ an untrained Convolutional Neural Network (CNN) to generate an image, which can serve as a sufficient image prior. This idea is rather similar to (Jeon et al., 2021) where a network is pre-trained for image generation. The final objective of the introduced attack method is
$\arg\min_{x}$ $\mathcal{L}_{\rm grad}(x;\theta,\nabla_{\theta}\mathcal{L}_{\theta}(x^{*},y^{*}))+\alpha_{TV}\mathcal{R}_{\rm TV}$
$+\alpha_{BN}\mathcal{R}_{\rm BN}+\alpha_{f}\mathcal{R}_{\rm feature}(x).$
5 Experiments
5.1 Experiment setup
Key parameters of defenses.
Key experimental setup for the attack.
We use a stratified sampled subset of 50 CIFAR-10 images to evaluate the attack performance. For all attack algorithms, we assume the batch norm statistics are available following (Geiping et al., 2020). We observe the difficulty of optimization when using cosine-similarity as the gradient inversion loss for the proposed attack algorithm and therefore address this issue by reweighting the gradients by their $\ell_{0}$ norm. See details in the Appendix FGeiping et al., 2020; Zhu et al., 2019 have shown that small batch size is vital for the success of the attack. We intentionally evaluate the attack with two small batch sizes $4$ and $8$ and two small but realistic batch sizes, 16 and 32.
Metrics for reconstruction quality.
5.2 Experiment results
@@
In Section 3, we analyze the performance of different defenses on 2-layer networks. In the experiments, our attack method involves intermediate feature matching, which requires reconstructing feature from a 2-layer network. This setting aligns with our theoretical analysis. Therefore, we can verify the theory by checking the feature reconstruction error in our attack. The results is shown in Table 2. The increase of the reconstruction error along with the batch size matches the theoretical bounds. Moreover, gradient pruning and gradient noise with large scaling has significantly large reconstruction error, which also aligns with our theoretical evaluation.
@@
According to the framework of evaluation, we need to consider strong attacks against various defenses. A key motivation of our proposed attack is its good performance among different defense, which we will verify in the following results. Our method is based on gradient inversion and can be easily added to previous methods (Geiping et al., 2020; Zhu et al., 2019). In Table  3, we compare the state-of-the-art gradient inversion methods with the proposed attacking method, our method outperforms previous methods. We visualize reconstructed images in Figure 2. Without defense, both (Geiping et al., 2020) and our attack method can recover images well from the gradient, and our method produces higher-quality images. With defenses, our method shows better robustness – the recovered images are visually more similar to the original image. See Table 7 and Table 8 in Appendix I for summary of reconstruction results.
@@
In our systematic evaluation of different defenses, we measure their strength with the strongest attack: $\mathcal{S}_{D}=\max_{A\in\mathcal{A}}d(S,A(D(G)))$. In order to estimate the maximum reconstruction error, we selected four different attacks with various parameters and batch sizes to conduct the experiment and record the reconstruction RMSE in Table 4. For each setting, we select the attack with the smallest RMSE as the strongest attack. Therefore, among all defenses, gradient pruning with large $p$ and gradient noise with large variance $\sigma_{0}^{2}$ has the best defensive effect.
In order to evaluate the defenses, another key criteria is the utility loss of the defense methods. Most of the defenses prevent models from data leakage by perturbing the gradients in training, which may hurt the training task itself. In our experiment, we measure the utility by the final loss of the training task and we show it in Table 5. Gradient pruning, as one of the defenses having strongest effect towards attacks, causes smaller utility loss than gradient noise. This result indicates that gradient pruning is a better defense than gradient noise. In Figure 3, we plot the relation between the reconstruction error and the utility loss, where each dot represents a method under the setting of a specific batch size and parameter of the defense. With the same level of utility loss, gradient pruning has the largest RMSE, indicating it is the best defense to have $\min_{D\in\mathcal{D}_{\mathcal{U}}}\mathcal{S}_{D}$.